# Jiwen Yu1∗†  Yiran Qin1∗   Xintao Wang2‡   Pengfei Wan2  Di Zhang2  Xihui Liu1‡

### v1

@@ Abstract
Generative game engines have the potential to revolutionize game development by autonomously creating new content and reducing manual workload. However, existing video-based game generation methods fail to address the critical challenge of scene generalization, limiting their applicability to existing games with fixed styles and scenes. In this paper, we present GameFactory, a framework focused on exploring scene generalization in game video generation. To enable the creation of entirely new and diverse games, we leverage pre-trained video diffusion models trained on open-domain video data. To bridge the domain gap between open-domain priors and small-scale game dataset, we propose a multi-phase training strategy that decouples game style learning from action control, preserving open-domain generalization while achieving action controllability. Using Minecraft as our data source, we release GF-Minecraft, a high-quality and diversity action-annotated video dataset for research. Furthermore, we extend our framework to enable autoregressive action-controllable game video generation, allowing the production of unlimited-length interactive game videos. Experimental results demonstrate that GameFactory effectively generates open-domain, diverse, and action-controllable game videos, representing a significant step forward in AI-driven game generation. Our dataset and project page are publicly available at https://vvictoryuki.github.io/gamefactory/.
@@ 1 Introduction
Video diffusion models have demonstrated impressive capabilities in video generation and real-world physics simulation [26, 24, 48, 45]. This physics simulation ability can be applied to generate game physics, making them promising candidates for game engines [36, 2, 11]. These AI-driven engines have the potential to transform game development by significantly reducing manual work in the traditional game industry through automatic creation of game content. Given this promising direction, it is important to explore how to build a qualified generative game engine.
Generative game engines are typically implemented as video generation models with action controllability, enabling responses to user inputs like keyboard and mouse interactions. Scene generalization, as a crucial yet understudied area, enables the creation of new game scenes beyond existing games. Current research works [36, 2, 11, 3, 42] mainly focus on specific games such as DOOM [36, 42], Atari [2], CS:GO [2], Super Mario Bros [42], and Minecraft [11], or uses limited game-specific datasets [3]. This game-specific approach restricts models to generating content only for existing games, limiting their potential for creating new games. Thus, scene generalization remains a key challenge in advancing generative game engines.
@@
To achieve scene generalization in game videos, collecting large-scale action-annotated video datasets would be the most direct approach. With sufficient data covering all possible scenarios, it could enable arbitrary game scene generation. However, such annotation is prohibitively expensive and impractical for open-domain scenarios. In contrast, open-domain videos without action annotations are abundant on the Internet and contain rich generative priors. Leveraging scene-generation priors trained on these large-scale open-domain datasets presents a more feasible path to scene generalization. As shown in Figure 2, using a video generation model pre-trained on large-scale open-domain video data, we can train an effective action control module with just a small amount of action-annotated data, achieving controllability while maintaining the powerful generative prior. We chose Minecraft [13] as our action data source for its unique advantages: diverse and complex action space with customizable inputs, frame-by-frame action annotation support, and cost-effective generation of high-quality action data without human bias.
We present GameFactory, a method for open-domain scene generalization in game video generation. GameFactory has several key designs. First, directly fine-tuning the pretrained model on Minecraft data not only imparts action control capabilities, but also risks embedding a distinct Minecraft style into the generated games, thereby compromising the model’s open-domain generalization. To address this issue, we propose a multiphase decoupled training strategy (Section 3.5) that separates the learning of the game style from the learning of action control. Second, we develop different control mechanisms specifically designed for the distinct nature of continuous mouse movements versus discrete keyboard inputs in the action control module (Section 3.3). Finally, we extend our model to achieve autoregressive long video generation (Section  3.6). Autoregressive generation is particularly desirable, as it enables the production of videos with unlimited length, better satisfying the practical requirements of game video generation.
@@
  • We propose GameFactory for open-domain scene generalization in game video generation. To overcome game-specific limitations, we leverage pre-trained video diffusion models with a multiphase training strategy that decouples game style from action control, enabling diverse interactive game environments beyond existing games.
  • We introduce GF-Minecraft, a new action-annotated dataset for generative game engines. To address the challenge of human bias in action annotation, we curate a high-quality and diverse dataset from Minecraft.
  • We develop an autoregressive generation mechanism that enables unlimited-length action-controllable game videos, meeting the critical requirements of continuous gameplay in practical game applications.
These advancements pave the way for AI-driven game engines with both scene generalization and practical utility.
@@ 2.3 Game Video Generation
Given the promising capabilities of video diffusion models, researchers have begun exploring their application in game generation [7, 2, 36, 3, 11, 14, 42, 12]. Genie [7] proposes a foundation model for playable world based on video generation. DIAMOND [2], GameNGen [36], Oasis [11] and PlayGen [42] leverage diffusion-based world modeling for specific games like Atari [2], CS:GO [2], DOOM [36, 42], Minecraft [11] and Super Mario Bro [42]. GameGenX [3] introduces OGameData for game video generation and control. However, these works, including some of our concurrent works [3, 11, 42], often suffer from overfitting to specific games or datasets, exhibiting limited scene generalization capabilities. Other concurrent works, such as Matrix [14] and Genie 2 [12], have discussed the issue of achieving control generalization. However, they still rely on collecting large amounts of high-cost action-labeled data to learn sufficient knowledge about game scenarios.
@@
In contrast, our proposed GameFactory addresses scene generalization by utilizing pretrained video model priors and the easily-accessible and low-cost game dataset GF-Minecraft, with experiments validating its effectiveness.
3 Method
Figure 3: (a) Integration of Action Control Module into transformer blocks of the video diffusion model. (b) Different control mechanisms for continuous mouse signals and discrete keyboard signals (detailed analysis in Sec 3.3). (c) Due to temporal compression (compression ratio $r=4$), the number of latent features differs from the number of actions, causing granularity mismatch during fusion. Grouping aligns these sequences for fusion. Additionally, the $i$-th latent feature can fuse with action groups within a previous window (window size $w=3$), accounting for delayed action effects (e.g., ‘jump’ key affects several subsequent frames).
The key to creating new games using video generation models lies in scene generalization. However, it is impractical to fit all possible game scenarios merely by collecting large-scale gaming video datasets, as action annotations are expensive and the datasets can never truly achieve complete open-domain coverage. To address this challenge, we propose leveraging the open-domain generation prior of pre-trained models, combined with a small amount of action-annotated dataset to achieve effective action control with strong scene generalization capabilities.
Sec. 3.1 introduces preliminaries. Sec. 3.2 describes action-annotated data collection. Sec. 3.3 presents the action control module. Sec. 3.4 analyzes the challenges in maintaining the pre-trained model’s open-domain prior while achieving action control. Sec. 3.5 presents a multi-phase training strategy as the solution. Sec. 3.6 describes autoregressive generation for long video synthesis, which is essential for game video applications.
3.1 Preliminaries
@@
We adopt a transformer-based latent video diffusion model [27, 26] as the backbone. Let $\mathbf{X}$ represent a video sequence. To reduce modeling complexity, an Encoder $E(\cdot)$ compresses the video spatially and temporally into a latent representation $\mathbf{Z}=E(\mathbf{X})$. With temporal compression ratio $r$, a $(1+rn)$-frame video is compressed into $(1+n)$ latent frames. Denoting the $i$-th frame as $\mathbf{x}^{i}$ and $i$-th latent as $\mathbf{z}^{i}$, we have $\mathbf{X}=[\mathbf{x}^{0},\mathbf{x}^{1},...,\mathbf{x}^{rn}]$ and $\mathbf{Z}=[\mathbf{z}^{0},\mathbf{z}^{1},...,\mathbf{z}^{n}]$. To generate videos, we first train a noise predictor in the latent video diffusion model. During training, noise is added to the clean latent $\mathbf{Z}_{0}$. Let $\mathbf{Z}_{t}$ denote the noisy latent at timestep $t$. The noise addition process is formulated as:
$\mathbf{Z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{Z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $\sqrt{\bar{\alpha}_{t}}$ and $\sqrt{1-\bar{\alpha}_{t}}$ are predefined scheduler parameters, and $\boldsymbol{\epsilon}$ is a random noise sampled from the standard normal distribution. The noise predictor is trained using the following loss function:
$\mathcal{L}(\boldsymbol{\theta})=\mathbb{E}[||\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\mathbf{Z}_{t},\mathbf{p},t)-\boldsymbol{\epsilon}||_{2}^{2}],$ (2)
where $\boldsymbol{\theta}$ denotes the model parameters and $\mathbf{p}$ is the prompt input. During inference, with the trained noise predictor, we can sample clean latent $\mathbf{Z}_{0}$ from a noisy latent $\mathbf{Z}_{T}$. The predicted latent $\mathbf{Z}_{0}$ is then decoded back to video $\mathbf{X}$ through $D(\cdot)$: $\mathbf{X}=D(\mathbf{Z}_{0})$. When considering action control, the action $\mathbf{A}=[\mathbf{a}^{1},\mathbf{a}^{2},...,\mathbf{a}^{rn}]$, where $\mathbf{a}^{i}$ represents the action taken at timestep $(i-1)$ to transfer from $\mathbf{x}^{i-1}$ to $\mathbf{x}^{i}$. The corresponding action-conditioned loss function is:
@@
where $\boldsymbol{\phi}$ means parameters of the action control module.
@@ 3.2 GF-Minecraft Dataset
For our action-controllable video generation model to simulate real game engines, the training data should satisfy three key requirements: (1) easily accessible with customizable action inputs to enable cost-effective data collection; (2) action sequences free from human bias, allowing extreme and low-probability action combinations to support arbitrary action inputs; (3) diverse game scenes with corresponding textual descriptions to learn scene-specific physical dynamics. As existing game video datasets fail to meet these criteria, we introduce our GF-Minecraft dataset, where ‘GF’ stands for our method GameFactory and ‘Minecraft’ refers to the game name. The advantages of our dataset are summarized as follows and its details can be found in the Appendix A.
Minecraft as an Accessible Data Source.
We leverage Minecraft as our data collection platform for its comprehensive API that captures detailed environmental snapshots, enabling large-scale data collection with action annotations. The game also offers extensive scenes, navigable areas, diverse action space, and an open-world environment. By executing predefined action sequences, we collected 70 hours of gameplay video as GF-Minecraft Dataset.
Collecting Videos with Unbiased Action.
Existing Minecraft datasets, such as VPT [4], are collected from real human gameplay, resulting in biased action distributions that favor common human behaviors. Models trained on these datasets overlook rare action combinations, such as moving backward, jumping in place, or standing still with mouse movements. To eliminate such biases, we decompose keyboard and mouse inputs into atomic actions and ensure their balanced distribution. We also randomize the frame duration of each atomic action to avoid temporal bias.
Diverse Scenes with Textual Descriptions.
To enhance dataset diversity, we captured videos across different scenes, weather conditions, and times of day. We segmented the videos and annotated them with textual descriptions using the pre-trained multimodal large language model MiniCPM [46].
@@ 3.3 Action Control Module
We incorporate action control modules into the transformer blocks of the video diffusion model to enable action-controllable generation. The architecture of the transformer block is shown in Fig. 3 (a), and the structure of the action control module is shown in Fig. 3 (b). Suppose that the input action includes both continuous mouse movement action $\mathbf{M}$ and discrete keyboard action $\mathbf{K}$. The intermediate feature in the Transformer is denoted as $\mathbf{F}$:
$\mathbf{M}$ $=[\mathbf{m}^{1},\mathbf{m}^{2},...,\mathbf{m}^{rn}]\in\mathbb{R}^{rn\times d_{1}},$ (4)
$\mathbf{K}$ $=[\mathbf{k}^{1},\mathbf{k}^{2},...,\mathbf{k}^{rn}]\in\mathbb{R}^{rn\times d_{2}},$ (5)
$\mathbf{F}$ $=[\mathbf{f}^{0},\mathbf{f}^{1},...,\mathbf{f}^{n}]\in\mathbb{R}^{(n+1)\times l\times c},$ (6)
where $\mathbf{m}^{i}$ represents the $i$-th mouse action, $\mathbf{k}^{i}$ represents the $i$-th keyboard action, $\mathbf{f}^{i}$ represents the feature of the $i$-th latent frame. The total number of video frames is $(1+rn)$, compressed to $(1+n)$ latent frames, where $r$ denotes the temporal compression ratio. $d_{1}$ and $d_{2}$ denote the dimension numbers of actions, $l$ is the length of token sequence and $c$ is the number of feature channels.
Grouping Actions with a Sliding Window.
@@
Due to the temporal compression ratio $r$, the number of actions ($rn$) differs from the number of features ($n+1$), creating a granularity mismatch for action-feature fusion. As shown in Fig. 3 (c), we address this by grouping actions using a sliding window of size $w$. For the $i$-th feature $\mathbf{f}^{i}$, we consider actions within $[\mathbf{a}^{r\times(i-w+1)},...,\mathbf{a}^{ri}]$. This window design captures delayed action effects, such as how a jump command influences multiple subsequent frames. For out-of-range indices, boundary actions are used as padding. For mouse movement actions $\mathbf{M}$, the grouping operation can be represented as:
$\mathbf{M}_{group}=\mathbf{group}(\mathbf{M})\in\mathbb{R}^{(n+1)\times rw\times d_{1}}.$ (7)
As for keyboard actions $\mathbf{K}$, we first learn the embedding of actions and add the positional encoding to get:
$\mathbf{K}_{embed}=\mathbf{embed}(\mathbf{K})\in\mathbb{R}^{rn\times c}.$ (8)
After that, perform a grouping operation on the action embeddings, which can be represented as follows:
$\mathbf{K}_{group}=\mathbf{group}(\mathbf{K}_{embed})\in\mathbb{R}^{(n+1)\times rw\times c}.$ (9)
Mouse Movements Control.
@@
To fuse the grouped mouse action $\mathbf{M}_{group}$ with feature $\mathbf{F}$, we first reshape it from $\mathbb{R}^{(n+1)\times rw\times d_{1}}$ to $\mathbb{R}^{(n+1)\times 1\times rwd_{1}}$. Then we repeat it in the dimension of the token sequence length to get:
$\mathbf{M}_{repeat}=\mathbf{repeat}(\mathbf{M}_{group})\in\mathbb{R}^{(n+1)\times l\times rwd_{1}},$ (10)
and concatenate it with $\mathbf{F}$ along with the channel dimension:
$\mathbf{F}_{fused}=\mathbf{cat}([\mathbf{F},\mathbf{M}_{repeat}])\in\mathbb{R}^{(n+1)\times l\times(c+rwd_{1})}.$ (11)
After that, further learning on $\mathbf{F}_{fused}$ is conducted through a layer of MLP and a layer of temporal self-attention.
Keyboard Actions Control.
Figure 4: Style bias in video game generation. The model tuned on Minecraft data inherits its distinctive pixelated block style, creating a domain gap from the original parameters. This motivates decoupling action control learning from data style learning through specialized training strategies.
@@
For discrete keyboard control, we perform a cross-attention calculation between the grouped action embeddings $\mathbf{K}_{group}\in\mathbb{R}^{(n+1)\times rw\times c}$ and $\mathbf{F}\in\mathbb{R}^{(n+1)\times l\times c}$, similar to the prompt cross-attention between text and $\mathbf{F}$. Specifically, $\mathbf{K}_{group}$ serves as the key and value in the attention, while $\mathbf{F}$ functions as the query.
Action-Specific Control Module.
Our experiments indicate concatenation works better for continuous actions (e.g., mouse movements) while cross-attention suits discrete actions (e.g., keyboard inputs). For continuous actions, concatenation preserves the magnitude information critical for control, whereas cross-attention’s similarity computations and normalization would diminish these numerical differences. For discrete actions, cross-attention’s effectiveness aligns with its successful applications in text prompt control, demonstrating its strength in categorical processing.
3.4 Leveraging Pre-trained Models for Action-Controlled Scene Generalization
Recent video generation models for games [2, 36, 11, 42], are limited to reproducing existing games rather than creating novel games, hindering the vision of a generative game engine that could reduce game development costs. While a large-scale open-domain game video dataset would be ideal, annotating such data is prohibitively expensive and achieving true open-domain coverage remains infeasible. We explore leveraging a pre-trained video generation model that can learn rich generative priors from unlabeled open-domain data. By fine-tuning an action control module using limited game data (e.g., our proposed GF-Minecraft Dataset) on this pre-trained model, we enable controlled game creation across open-domain scenes.
However, this approach faces a key challenge: achieving scene generalization without stylistic bias. While our model enables new game scene generation, the outputs tend to inherit the visual style of the training data (e.g., Minecraft’s pixelated blocks, Fig. 4). This style bias stems from the diffusion model’s loss function (Eq. 3), where learning action control naturally captures dataset-specific visual features that help minimize training loss. To address this, we next propose a multi-phase training strategy that decouples action control learning from visual style, enabling robust action control independent of specific data styles.
3.5 Multi-Phase Training Strategy
Phase #1: Tune LoRA to Fit Game Videos.
We fine-tune the pre-trained video diffusion model using LoRA [21] to adapt it to specific game video while preserving most original parameters. This produces a model specialized for the target game domain. Better style adaptation in this phase allows the next phase to focus purely on action control, reducing style-control entanglement.
Phase #2: Tune Action Control Module.
We freeze both pre-trained parameters and LoRA, only training the action control module with game videos and action signals. Since Phase #1 has handled style adaptation through LoRA, the training loss now focuses less on style learning. This allows the model to concentrate on action control learning, as it becomes the main contributor to minimizing the diffusion loss. Such separation enables style-independent control that can generalize to open-domain scenarios.
Phase #3: Inference on Open Domain.
During inference, we remove the LoRA weights for game style adaptation, keeping only the action control module parameters. Thanks to the decoupling in previous phases, the action control module can now work independently of specific game styles, enabling controlled game video generation across open-domain scenarios.

### v2

@@ Abstract
Generative videos have the potential to revolutionize game development by autonomously creating new content. In this paper, we present GameFactory, a framework for action-controlled scene-generalizable game video generation. We first address the fundamental challenge of action controllability by introducing GF-Minecraft, a action-annotated game video dataset without human bias, and developing a action control module that enables precise control over both keyboard and mouse inputs. We further extend to support autoregressive generation for unlimited-length interactive videos. More importantly, GameFactory tackles the critical challenge of scene-generalizable action control, which most existing methods fail to address. To enable the creation of entirely new and diverse games beyond fixed styles and scenes, we leverage the open-domain generative priors from pre-trained video diffusion models. To bridge the domain gap between open-domain priors and small-scale game datasets, we propose a multi-phase training strategy with a domain adapter that decouples game style learning from action control. This decoupling ensures that action control learning is no longer bound to specific game styles, thereby achieving scene-generalizable action control. Experimental results demonstrate that GameFactory effectively generates open-domain action-controllable game videos, representing a significant step forward in AI-driven game generation. Our dataset and project page are publicly available at https://yujiwen.github.io/gamefactory/.
@@ 1 Introduction
Video diffusion models have demonstrated impressive capabilities in video generation [31, 26, 52, 50]. These models have the potential to become promising candidates for generative game engines [42, 2, 11, 6, 12], which could transform game development by not only significantly reducing manual work in the traditional game industry through automatic content creation, but also enabling the generation of unlimited game content for players to explore. Given this promising direction, it is important to explore how to build a qualified generative game engine.
Generative game engines are typically implemented as video generation models with action controllability, enabling responses to user inputs like keyboard and mouse interactions [42, 11, 12]. Current research works [42, 2, 11, 7, 47] mainly focus on specific games such as DOOM [42, 47], Atari [2], CS:GO [2], Super Mario Bros [47], and Minecraft [11], or use limited game-specific datasets [7]. This game-specific approach lacks scene generalization capability, preventing models from creating content beyond existing games and limiting their potential for developing new games. Thus, scene generalization remains a key challenge in advancing generative game engines.
@@
To achieve scene generalization in game videos, collecting large-scale action-annotated video datasets would be the most direct approach. With sufficient data covering all possible scenarios, it could enable arbitrary game scene generation. However, such action annotations are prohibitively expensive and impractical for open-domain videos. In contrast, open-domain videos are abundant on the Internet, and video generation models trained on these videos contain rich generative priors. Leveraging these scene-generation priors presents a more feasible path to scene generalization (upper blue section in Fig. 2). This observation motivates us to ask: With the video generation models pre-trained on large-scale open-domain videos, along with small-scale action-annotated game videos, can we empower scene-generalizable video generation with action control? This is a nontrivial problem, because directly fine-tuning the pre-trained video generation models with action-annotated game videos will lead to degraded scene generalizability (i.e., the model will collapse to the specific domain of the game videos).
In this work, we present GameFactory, a framework for action-controlled and scene-generalizable game video generation. To achieve action-controlled video generation, we firstly collect an action-annotated game video dataset GF-Minecraft. This dataset, sourced from the Minecraft platform [13], features an unbiased action distribution, diverse scenes, and text annotations (Sec. 4.1). Secondly, we design action control mechanisms for continuous mouse movements and discrete keyboard inputs, respectively. The action control modules can be seamlessly injected into the pre-trained video diffusion models without affecting the video generation abilities of the pre-trained models (Sec. 4.2). Lastly, we extend our model to autoregressive long video generation by allowing varying noise levels across different frames, continuously conditioning on previously generated frames to generate new video frames (Sec. 4.3).
However, directly fine-tuning the pretrained model on Minecraft data not only imparts action control capabilities, but also risks embedding a distinct Minecraft style into the generated games, thereby compromising the model’s open-domain generalization. To address this challenge, our key idea is to disentangle the learning of game style and action control with different modules and parameters, so that we can remove the game style without affecting the action controllability for generating open-domain videos with action control (Sec. 5.1). Specifically, we propose a domain adapter to fit Minecraft style, combined with a multiphase decoupled training strategy (Sec. 5.2).
@@
  • We propose GameFactory for action-controlled and scene-generalizable video generation, enabling diverse interactive game generation beyond existing games.
  • We introduce GF-Minecraft, an action-annotated game video dataset featuring unbiased action distributions, diverse scenes, and text descriptions. We also design dedicated mechanisms for action control and autoregressive long video generation.
  • We propose a domain adapter with multiphase decoupled training strategy that disentangles game style from action control, enabling open-domain video generation while preserving action controllability.
@@ 2.3 Game Video Generation
Early works [25, 24, 27, 28, 29] began exploring game generation using generative models like GAN, but were primarily limited by the generative capability of these models. Recently, given the promising capabilities of video generation models, researchers have begun exploring their application in game generation [6, 2, 42, 7, 11, 14, 47, 12, 23]. Genie [6] proposes a foundation model for playable world based on video generation. DIAMOND [2], GameNGen [42], Oasis [11] and PlayGen [47] leverage diffusion-based world modeling for specific games like Atari [2], CS:GO [2], DOOM [42, 47], Minecraft [11] and Super Mario Bro [47]. GameGenX [7] introduces OGameData for game video generation and control. However, these works suffer from overfitting to specific games or datasets, exhibiting limited scene generalization capabilities. Recent works like Genie 2 [12] and Matrix [14] have made valuable progress towards control generalization in game video generation. While Genie 2 demonstrates impressive results through extensive action-labeled data collection, and Matrix shows promising generalization capabilities in racing games, there remains room for improvement in achieving broader scene generalization with more diverse action controls.
@@
In contrast, our proposed GameFactory addresses scene generalization by utilizing pretrained video model priors and the easily-accessible and low-cost game dataset GF-Minecraft, with experiments validating its effectiveness across diverse open-world scenarios and a much more complex action space (including forward/backward/left/right movement, jumping, and acceleration/deceleration).
3 Preliminaries
@@
We adopt a transformer-based latent video diffusion model [33, 31] as the backbone. Let $\mathbf{X}$ represent a video sequence. To reduce modeling complexity, an Encoder $E(\cdot)$ compresses the video spatially and temporally into a latent representation $\mathbf{Z}=E(\mathbf{X})$. With temporal compression ratio $r$, a $(1+rn)$-frame video is compressed into $(1+n)$ latent frames. Denoting the $i$-th frame as $\mathbf{x}^{i}$ and $i$-th latent as $\mathbf{z}^{i}$, we have $\mathbf{X}=[\mathbf{x}^{0},\mathbf{x}^{1},...,\mathbf{x}^{rn}]$ and $\mathbf{Z}=[\mathbf{z}^{0},\mathbf{z}^{1},...,\mathbf{z}^{n}]$. During training, noise is added to the clean latent $\mathbf{Z}_{0}$ to get noisy latent $\mathbf{Z}_{t}$ at timestep $t$. $\boldsymbol{\epsilon}$ is the added random noise. When considering action control, the action $\mathbf{A}=[\mathbf{a}^{1},\mathbf{a}^{2},...,\mathbf{a}^{rn}]$, where $\mathbf{a}^{i}$ represents the action taken at timestep $(i-1)$ to transfer from $\mathbf{x}^{i-1}$ to $\mathbf{x}^{i}$. The corresponding action-conditioned loss function is:
@@
where $\boldsymbol{\phi}$ means the model parameters, and $\mathbf{p}$ is the prompt input. During inference, we can sample clean latent $\mathbf{Z}_{0}$ from a noisy latent $\mathbf{Z}_{T}$. The predicted latent $\mathbf{Z}_{0}$ is then decoded back to video $\mathbf{X}$ through $D(\cdot)$: $\mathbf{X}=D(\mathbf{Z}_{0})$.
4 Action-Controlled Video Generation
This section introduces implementing an action-controlled video generation model, which is the foundation for Sec. 5. Sec. 4.1 describes our collected game video dataset and its advantages; Sec. 4.2 introduces how to implement the action control module for responding to player actions; Sec. 4.3 presents the method for autoregressive long video generation, which is the key to creating playable game.
@@ 4.1 GF-Minecraft Dataset
For our action-controllable video generation model to simulate real game engines, the training data should satisfy three key requirements: (1) easily accessible with customizable action inputs to enable cost-effective data collection; (2) action sequences free from human bias, allowing extreme and low-probability action combinations to support arbitrary action inputs; (3) diverse game scenes with corresponding textual descriptions to learn scene-specific physical dynamics. Existing datasets like VPT [3] are collected from human gameplay videos, which inherently contain human behavioral biases and lack text descriptions of the scenes. To address these limitations, we introduce our GF-Minecraft dataset, where ‘GF’ stands for our method GameFactory and ‘Minecraft’ refers to the game name. The advantages of our dataset are summarized as follows and its details can be found in the supplementary materials:
❑ Minecraft as an Accessible Data Source. We leverage Minecraft as our data collection platform for its comprehensive API that captures detailed environmental snapshots, enabling large-scale data collection with action annotations. The game also offers extensive scenes, navigable areas, diverse action space, and an open-world environment. By executing predefined action sequences, we collected 70 hours of gameplay video as GF-Minecraft Dataset.
❑ Collecting Videos with Unbiased Action. Existing Minecraft datasets, such as VPT [3], are collected from real human gameplay, resulting in biased action distributions that favor common human behaviors. Models trained on these datasets overlook rare action combinations, such as moving backward, jumping in place, or standing still with mouse movements. To eliminate such biases, we decompose keyboard and mouse inputs into atomic actions and ensure their balanced distribution. We also randomize the frame duration of each atomic action to avoid temporal bias.
❑ Diverse Scenes with Textual Descriptions. To enhance dataset diversity, we captured videos across different scenes, weather conditions, and times of day. We segmented the videos and annotated them with textual descriptions using an efficient multimodal LLM, MiniCPM [51].
@@ 4.2 Action Control Module
Figure 3: (a) Integration of Action Control Module into transformer blocks of the video diffusion model. (b) Different control mechanisms for continuous mouse and discrete keyboard inputs.
Figure 4: Due to temporal compression (compression ratio $r=4$), the number of latent features differs from the number of actions, causing granularity mismatch during fusion. Grouping aligns these sequences for fusion. Additionally, the $i$-th latent feature can fuse with action groups within a previous window (window size $w=3$), accounting for delayed action effects (e.g., ‘jump’ key affects several subsequent frames).
@@
We incorporate action control modules into the transformer blocks of the video diffusion model to enable action-controllable generation. The architecture of the transformer block is shown in Fig. 3 (a), and the structure of the action control module is shown in Fig. 3 (b). Suppose that the input action includes both continuous mouse movement action $\mathbf{M}\in\mathbb{R}^{rn\times d_{1}}$, and discrete keyboard action $\mathbf{K}\in\mathbb{R}^{rn\times d_{2}}$. The intermediate feature in the Transformer is denoted as $\mathbf{F}\in\mathbb{R}^{(n+1)\times l\times c}$. The total number of video frames is $(1+rn)$, compressed to $(1+n)$ latent frames, where $r$ denotes the temporal compression ratio. $d_{1}$ and $d_{2}$ denote the dimension numbers of actions, $l$ is the length of token sequence and $c$ is the number of feature channels.
@@
❑ Grouping Actions with a Sliding Window. Due to the temporal compression ratio $r$, the number of actions ($rn$) differs from the number of features ($n+1$), creating a granularity mismatch for action-feature fusion. As shown in Fig. 4, we address this by grouping actions using a sliding window of size $w$. For the $i$-th feature $\mathbf{f}^{i}$, we consider actions within $[\mathbf{a}^{r\times(i-w+1)},...,\mathbf{a}^{ri}]$. This window design captures delayed action effects, such as how a jump command influences multiple subsequent frames. For out-of-range indices, boundary actions are used as padding. For mouse movement $\mathbf{M}$, the grouped action is $\mathbf{M}_{group}\in\mathbb{R}^{(n+1)\times rw\times d_{1}}$. As for keyboard actions $\mathbf{K}$, we first learn the embedding of actions and add the positional encoding. After that, we perform a grouping operation on the action embeddings to get $\mathbf{K}_{group}\in\mathbb{R}^{(n+1)\times rw\times c}$.
@@
❑ Mouse Movements Control. To fuse the grouped mouse action $\mathbf{M}_{group}$ with feature $\mathbf{F}$, we first reshape it from $\mathbb{R}^{(n+1)\times rw\times d_{1}}$ to $\mathbb{R}^{(n+1)\times 1\times rwd_{1}}$. Then we repeat it in the dimension of the token sequence length to get $\mathbf{M}_{repeat}\in\mathbb{R}^{(n+1)\times l\times rwd_{1}}$, and concatenate it with $\mathbf{F}$ along with the channel dimension to get $\mathbf{F}_{fused}\in\mathbb{R}^{(n+1)\times l\times(c+rwd_{1})}$. After that, further learning on $\mathbf{F}_{fused}$ is conducted through a layer of MLP and a layer of temporal self-attention.
@@
❑ Keyboard Actions Control. For discrete keyboard control, we perform a cross-attention calculation between the grouped action embeddings $\mathbf{K}_{group}\in\mathbb{R}^{(n+1)\times rw\times c}$ and $\mathbf{F}\in\mathbb{R}^{(n+1)\times l\times c}$, similar to the prompt cross-attention between text and $\mathbf{F}$. Specifically, $\mathbf{K}_{group}$ serves as the key and value in the attention, while $\mathbf{F}$ functions as the query.